# Training-Free Bayesianization for Low-Rank Adapters of Large Language Models

### v1

@@ Abstract
Estimating the uncertainty of responses of Large Language Models (LLMs) remains a critical challenge. While recent Bayesian methods have demonstrated effectiveness in quantifying uncertainty through low-rank weight updates, they typically require complex fine-tuning or post-training procedures. In this paper, we propose Training-Free Bayesianization (TFB), a novel framework that transforms existing off-the-shelf trained LoRA adapters into Bayesian ones without additional training. TFB systematically searches for the maximally acceptable level of variance in the weight posterior, constrained within a family of low-rank isotropic Gaussian distributions. We theoretically demonstrate that under mild conditions, this search process is equivalent to variational inference for the weights. Through comprehensive experiments, we show that TFB achieves superior uncertainty estimation and generalization compared to existing methods while eliminating the need for complex training procedures. Code will be available at https://github.com/Wang-ML-Lab/bayesian-peft.
Machine Learning, ICML
@@ 1 Introduction
Despite recent advances in Large Language Models (LLMs) showing great capacity for generating responsive answers to human instructions (Biderman et al., 2023; Wei et al., 2022, 2021; Min et al., 2022; Chowdhery et al., 2023; Anil et al., 2023; Touvron et al., 2023a, b; Radford et al., 2019; Brown et al., 2020; Achiam et al., 2023; OpenAI, 2022), the reliability of such large models remains a critical concern for researchers (Wang et al., 2024b, a), as untruthful yet confident answers could cause significant damage to individuals and society (Gupta et al., 2024; Nikitin et al., 2024; Yadkori et al., 2024; Kapoor et al., 2024). The accurate estimation of uncertainty in LLMs has thus emerged as a crucial challenge. Current research approaches mainly follow two paths: one focuses on eliciting models’ internal (verbalized) uncertainty through direct questioning (Xiong et al., 2023; Tian et al., 2023; Kapoor et al., 2024), while the other employs fine-tuning (Kapoor et al., 2024; Wang et al., 2024d) or post-training procedures (Yang et al., 2023).
@@
Both approaches face inherent limitations. Verbalized uncertainty, while simple to implement, remains controversial in terms of its empirical reliability and theoretical soundness (Kadavath et al., 2022; Kuhn et al., 2023). On the other hand, low-rank adapters (LoRA (Hu et al., 2022)), which offer a parameter-efficient way to adapt LLMs by adding a small set of low-rank weight matrices, have emerged as a promising direction for fine-tuning models. However, while LoRA efficiently adapts large models to new tasks, it does not itself provide a mechanism for principled uncertainty estimation. Recent Bayesianization attempts, such as BLoB (Wang et al., 2024d), integrate Bayesian methods with LoRA but still require complex training procedures and delicate hyperparameter tuning, limiting their practicality. These constraints motivate the following research question:
Can we “Bayesianize” a low-rank adapter in a theoretically sound and empirically simple way?
@@
In this paper, leveraging the existing abundance of trained LLM adapters across various domains, we diverge from conventional fine-tuning and post-training approaches. Instead, we develop a Training-Free Bayesianization (TFB) technique applicable to any given low-rank LLM adapter. TFB constrains the family of full-weight posteriors produced by LoRA adapters to low-rank isotropic Gaussian distributions. Given a trained LoRA adapter, it systematically searches for the maximally acceptable variance of the variational distribution of the weight posterior. TFB’s search range and stopping criteria can be determined using any additional in-distribution “anchor dataset,” independent of its supervision status or prior use in LoRA training.
@@
Despite its straightforward approach, we theoretically demonstrate that under mild conditions, TFB’s process of finding the maximal variance of the isotropic Gaussian posterior is equivalent to variational inference given an isotropic Gaussian prior with sufficiently large variance.
@@
Through extensive empirical evaluation across various settings, datasets, LLM backbones, LoRA weights, and LoRA variants, we verify TFB’s effectiveness. Our comprehensive experiments demonstrate that this novel training-free Bayesianization framework consistently achieves superior generalization and more accurate uncertainty quantification. To summarize, the main contributions of this paper are:
@@
  • We propose Training-Free Bayesianization (TFB), the first framework to transform trained LoRA adapters into Bayesian ones without re-training, continued training, or even gradient estimation.
@@
  • We establish theoretical connections between TFB and Variational Inference (VI), proving the equivalence between our variance maximization and variational free energy minimization under mild conditions.
@@
  • We develop an efficient, practical implementation requiring only an anchor dataset for search, making it widely applicable across different domains and compatible with various LoRA variants.
@@
  • Through comprehensive experiments, we demonstrate TFB consistently improves uncertainty estimation for off-the-shelf LoRA adapters, and overall surpasses the state-of-the-art counterparts of Bayesian LoRA.
Notation. Scalars, vectors, and matrices are denoted by lowercase letters, lowercase boldface letters, and uppercase boldface letters, respectively. For a matrix ${\bm{X}}=[{\bm{x}}_{1},\cdots,{\bm{x}}_{n}]\in\mathbb{R}^{m\times n}$, we use $\operatorname{vec}({\bm{X}})=[{\bm{x}}_{1}^{\top},{\bm{x}}_{2}^{\top},\cdots,{\bm{x}}_{n}^{\top}]^{\top}\in\mathbb{R}^{(mn)\times 1}$ to denote vectorization. $\otimes$ and $\circ$ are used to denote the Kronecker and element-wise product, respectively.
@@ 2 Related Work
LLM Uncertainty Estimation. To estimate the uncertainty of large language models (LLMs), the models themselves are often employed to generate and evaluate their own uncertainty (Lin et al., 2022; Kadavath et al., 2022). However, such approaches typically rely on task-specific labels and require additional training. Semantic entropy (Kuhn et al., 2023) leverages the invariance of language stemming from shared meanings to estimate uncertainty, while mutual information is used to compute a lower bound on model uncertainty by sampling from the model’s output distribution (Yadkori et al., 2024). Despite their contributions, these methods fail to accurately capture true model uncertainty, as they do not model the probability distribution over the parameter space (Hüllermeier & Waegeman, 2021; Abdar et al., 2021; Gawlikowski et al., 2023).
@@
Bayesian Low-Rank Adaptation. The Bayesian framework provides a powerful approach for capturing and estimating uncertainty by defining prior distributions and approximating posterior distributions over the parameter space (Neal, 2012; Hernández-Lobato & Adams, 2015; Gal & Ghahramani, 2016; Wang & Yeung, 2016; Gustafsson et al., 2020). However, modeling parameter distributions across the entire parameter space during fine-tuning introduces significant computational overhead (Fan et al., 2020; Zhang et al., 2021). To address this challenge, recent research has explored combining Bayesian methods with Parameter-Efficient Fine-Tuning (PEFT) techniques to improve the efficiency of uncertainty estimation. Several notable approaches have emerged in this direction. Wang et al. (2023) and Balabanov & Linander (2024) demonstrate improved performance by training multiple LoRA modules and ensemble their predictions during inference. Taking a different approach, Yang et al. (2023) applies a Kronecker-factorized Laplace approximation to fine-tuned LoRA parameters. More recently, BLoB (Wang et al., 2024d) advances the field by simultaneously estimating both the mean and covariance of LLM parameters within a single fine-tuning stage, leading to substantial performance improvements. Our proposed training-free Bayesianization represents a significant departure from these existing methods. Unlike approaches that require re-training (Gal & Ghahramani, 2016; Wang et al., 2023; Balabanov & Linander, 2024; Wang et al., 2024d) or rely on continued training and gradient estimation (Yang et al., 2023), our method achieves uncertainty estimation without any additional training steps, substantially improving the simplicity and efficiency for Bayesian learning of LLMs.
3 Methodology
@@
In this section, we present Training-Free Bayesianization (TFB), detailing our approach across four subsections. Sec. 3.1 establishes the foundational problem setting and preliminaries. Sec. 3.2 introduces our low-rank isotropic Gaussian posterior formulation, while Sec. 3.3 presents our core contribution: a novel approach for converting deterministic weights to probabilistic distributions without training. Sec. 3.4 provides the complete algorithmic implementation. The theoretical foundations supporting our method are separately addressed in Sec. 4.
@@ 3.1 Preliminaries
Low-Rank Adapation (LoRA). Given a pre-trained neural network layer with weight matrix ${\bm{W}}_{0}$, Low-Rank Adaptation (LoRA) (Hu et al., 2022) confines weight updates to a low-rank subspace during fine-tuning, expressing the update as $\Delta{\bm{W}}={\bm{B}}{\bm{A}}$, where $\Delta{\bm{W}}\in\mathbb{R}^{m\times n}$, ${\bm{B}}\in\mathbb{R}^{m\times r}$, and ${\bm{A}}\in\mathbb{R}^{r\times n}$. For input ${\bm{h}}$ and output ${\bm{z}}$ of the LoRA layer, the forward pass computation is given by:
@@
LoRA Bayesianization with Degenerate Gaussian. BLoB (Wang et al., 2024d), a pioneering work in low-rank Bayesianization for LLMs, empirically demonstrates that applying Gaussian distributions independently to elements of ${\bm{A}}$ suffices for effective uncertainty estimation in LoRA. Specifically, each element follows $q(A_{ij})={\mathcal{N}}(A_{ij}|M_{ij},\Omega_{ij}^{2}),\forall i\in[r],\forall j\in[n]$, where matrices ${\bm{M}}$ and ${\bm{\Omega}}$, sharing the dimensions of ${\bm{A}}$, represent the mean and standard deviation of the random variable ${\bm{A}}$, respectively. This formulation yields a low-rank degenerate structure in the full-weight space for the Bayesianized low-rank adapter’s Gaussian posterior:
@@
where ${\bm{\mu}}_{q}=\operatorname{vec}({\bm{W}}_{0}+{\bm{B}}{\bm{M}})$ denotes the mean, and ${\bm{\Sigma}}_{q}=[{\bm{I}}_{n}\otimes{\bm{B}}][\operatorname{diag}(\operatorname{vec}({\bm{\Omega}})^{2})][{\bm{I}}_{n}\otimes{\bm{B}}^{\top}]$ represents the low-rank degenerate covariance matrix.
In this work, we adopt a similar approach for modeling the variational distribution of the weight posterior by employing a Gaussian distribution with a low-rank degenerate covariance matrix, focusing exclusively on Bayesianizing the weight update matrix ${\bm{A}}$.
3.2 Low-Rank Isotropic Gaussian Posterior
@@
where ${\bm{\mu}}_{q}$ is defined as in Eqn. 2. Here, $\sigma_{q}^{2}{\bm{I}}\in\mathbb{R}^{mn\times mn}$ represents a full-rank isotropic covariance matrix with standard deviation $\sigma_{q}$, and $\operatorname{proj}$ denotes a linear projection operator that maps the full-space covariance matrix onto the low-rank space jointly defined by ${\bm{B}}$ and ${\bm{A}}$.
@@
The choice of low-rank isotropic Gaussian posteriors serves both theoretical and empirical purposes: it provides a single-parameter family that enables converting the VI objective into a maximization problem (Sec. 3.3 and Theorem 4.2), and empirically outperforms alternative distribution families (Sec. 5.3). We now present a practically efficient implementation for Bayesianizing LoRA under the constraint specified in Eqn. 3, with detailed theoretical analysis provided in Theorem 4.1.
@@
TFB in Practice. Consider a LoRA layer with weight updates ${\bm{B}}\in\mathbb{R}^{m\times r},{\bm{A}}\in\mathbb{R}^{r\times n}$ and a standard deviation scale $\sigma_{q}>0$. We begin by computing the compact Singular Value Decomposition (SVD) (Klema & Laub, 1980) of ${\bm{B}}$:
${\bm{B}}$ $={\bm{U}}{\bm{D}}{\bm{V}}^{\top},$ (4)
@@
where ${\bm{U}}\in\mathbb{R}^{m\times r}$ and ${\bm{V}}\in\mathbb{R}^{r\times r}$ are orthonormal matrices, and ${\bm{D}}=\operatorname{diag}({\bm{d}})\in\mathbb{R}^{r\times r}$ is a diagonal matrix with positive entries111 By stating ${\bm{d}}\succ{\bm{0}}$, we assume the original ${\bm{B}}$ has the full column rank $r$, which is usually the case for LLM adaptation (even for random initialization). . We then transform the original weight matrices $\{{\bm{B}},{\bm{A}}\}$ into an equivalent pair
$\{{\bm{B}}^{\prime}$ $={\bm{U}}{\bm{D}},{\bm{A}}^{\prime}={\bm{V}}^{\top}{\bm{A}}\},$ (5)
@@
maintaining the equality $\Delta{\bm{W}}={\bm{B}}{\bm{A}}={\bm{B}}^{\prime}{\bm{A}}^{\prime}$. Following BLoB’s Asymmetric Bayesianization scheme, we define the variational distribution for ${\bm{A}}^{\prime}$ using mean matrix ${\bm{M}}={\bm{A}}^{\prime}$ and standard deviation matrix ${\bm{\Omega}}\in\mathbb{R}^{r\times n}$, such that $q(A^{\prime}_{ij})={\mathcal{N}}(A^{\prime}_{ij}|M_{ij},\Omega_{ij}^{2}),\forall i\in[r],\forall j\in[n]$. Unlike BLoB, our ${\bm{\Omega}}$ is not freely parameterized but instead derived from projecting the full-space matrix $\sigma_{q}{\bm{I}}$ onto the low-rank weight space:
$\Omega_{ij}$ $=\nicefrac{{\sigma_{q}}}{{d_{i}}},\quad\forall i\in[r],\forall j\in[n].$ (6)
@@
This solution can be expressed compactly as ${\bm{\Omega}}=[\nicefrac{{1}}{{{\bm{d}}}},\cdots,\nicefrac{{1}}{{{\bm{d}}}}]$, comprising $n$ repeated vectors. This formulation significantly improves memory efficiency during inference, reducing the storage requirements for standard deviation parameters from $O(rn)$ to $O(r)$. While alternative parameterization approaches are possible, they must be capable of generating the low-rank isotropic Gaussian noises as demonstrated in Theorem 4.1. We have selected the current method to ensure maximum compatibility with existing implementations (Wang et al., 2024d). In TFB, we maintain a single shared $\sigma_{q}$ across all LoRA layers.
3.3 Training-Free Bayesianization (TFB)

### v2

@@ Abstract
Estimating the uncertainty of responses from Large Language Models (LLMs) remains a critical challenge. While recent Bayesian methods have demonstrated effectiveness in quantifying uncertainty through low-rank weight updates, they typically require complex fine-tuning or post-training procedures. In this paper, we propose Training-Free Bayesianization (TFB), a simple yet theoretically grounded framework that efficiently transforms trained low-rank adapters into Bayesian ones without additional training. TFB systematically searches for the maximally acceptable level of variance in the weight posterior, constrained within a family of low-rank isotropic Gaussian distributions. Our theoretical analysis shows that under mild conditions, this search process is equivalent to KL-regularized variational optimization, a generalized form of variational inference. Through comprehensive experiments, we show that TFB achieves superior uncertainty estimation and generalization compared to existing methods while eliminating the need for complex Bayesianization training procedures. Code will be available at https://github.com/Wang-ML-Lab/bayesian-peft.
@@ 1 Introduction
Despite recent advances in Large Language Models (LLMs) showing great capacity for generating responsive answers to human instructions [5, 63, 62, 44, 9, 3, 53, 54, 49, 8, 2, 48], the reliability of such large models remains a critical concern [58, 57], as untruthful yet confident answers could cause significant damage to individuals and society [17, 47, 65, 28]. The accurate estimation of uncertainty in LLMs has thus emerged as an urgent challenge. Current approaches mainly follow two paths: one focuses on directly asking the model to elicit its internal internal (verbalized) uncertainty [64, 51, 28], while the other employs complex fine-tuning techniques [28, 66, 61].
@@
Both approaches suffer from inherent limitations. Verbalized uncertainty, while simple to implement, remains controversial in terms of its empirical reliability and theoretical soundness [27, 38]. On the other hand, low-rank adapters (LoRA [23]), which offer a parameter-efficient way to adapt LLMs by adding a small set of low-rank weight matrices, have emerged as a promising direction for fine-tuning models. However, while LoRA efficiently adapts large models to new tasks, it does not itself provide a mechanism for principled uncertainty estimation. In response, recent Bayesianization attempts [66, 61], integrate Bayesian methods with LoRA, but they still require complex training procedures and sophisticated hyperparameter tuning, limiting their practicality. These constraints motivate the following research question:
Can we “Bayesianize” LLM low-rank adapters in a theoretically sound yet empirically simple way?
@@
In this paper, we diverge from conventional fine-tuning and post-training approaches. Instead, we develop a Training-Free Bayesianization (TFB) technique applicable to any given low-rank LLM adapter. TFB constrains the family of full-weight approximate posteriors produced by LoRA adapters to low-rank isotropic Gaussian distributions. Given a trained LoRA adapter, it systematically searches for the maximally acceptable variance of the variational distribution of the weight posterior, without the need for complex fine-tuning procedures. TFB’s search range and stopping criteria can be determined using any in-distribution “anchor dataset,” e.g., a small subset of the training dataset. Note that (1) this eliminates the need for an additional calibration or validation dataset; (2) this flexibility extends to both supervised and unsupervised data, even regardless of whether it was used in the original LoRA training.
@@
Despite its simplicity, we theoretically demonstrate that, TFB’s process of finding the maximal variance of the low-rank isotropic Gaussian posterior is equivalent to generalized variational inference, under mild conditions.
@@
We verify TFB’s effectiveness through extensive empirical evaluation across various settings, datasets, LLM backbones, LoRA weights, and LoRA variants. Our comprehensive experiments demonstrate that this novel training-free Bayesianization framework consistently achieves superior generalization and more accurate uncertainty estimation. To summarize, the main contributions of this paper are:
@@
  • We propose Training-Free Bayesianization (TFB), the first framework to transform trained LoRAs into Bayesian ones without re-training, continued training, or gradient estimation.
@@
  • We establish theoretical connections between TFB and generalized variational inference, proving their equivalence under mild conditions.
@@
  • We develop an efficient implementation of TFB requiring only an anchor dataset for search, making it widely applicable across different application scenarios.
@@
  • Through comprehensive experiments, we demonstrate that TFB consistently improves uncertainty estimation for off-the-shelf LoRA adapters, and overall surpasses the state-of-the-art counterparts of Bayesian LoRA.
@@ 2 Related Work
LLM Uncertainty Estimation. To estimate the uncertainty of LLMs, the models are often employed to generate and evaluate their own uncertainty [40, 27]. However, such approaches typically rely on task-specific labels and require additional training. Semantic entropy [38] leverages the invariance of language stemming from shared meanings to estimate uncertainty, while mutual information is used to compute a lower bound on model uncertainty by sampling from the model’s output distribution [65]. Despite their contributions, these methods fail to accurately capture true model uncertainty, as they do not model the probability distribution over the LLM parameters [25, 1, 15].
@@
Bayesian Low-Rank Adaptation. The Bayesian framework provides a powerful approach for capturing and estimating uncertainty during fine-tuning by defining prior distributions and approximating posterior distributions over the model parameters [46, 21, 14, 55]. Recent research has explored combining Bayesian methods with LoRA to mitigate the additional computational overhead associated with modeling parameter distributions across the entire parameter space. Yang et al. [66] applies a Kronecker-factorized Laplace approximation to fine-tuned LoRA parameters. More recently, BLoB [61] advances the field by simultaneously estimating both the mean and covariance of LLM parameters within a single fine-tuning stage. Our proposed training-free Bayesianization represents a significant departure from these existing methods. Unlike approaches that require re-training [14, 60, 4, 61] or rely on continued training and gradient estimation [66], our method achieves uncertainty estimation without any additional training steps, substantially improving the simplicity and efficiency for Bayesian learning of LLMs.
3 Training-Free Bayesianization (TFB)
@@
This section introduces our Training-Free Bayesianization (TFB). Sec. 3.1 introduces the problem setup. Sec. 3.2 and Sec. 3.3 present the two key parts of TFB: low-rank Gaussian variational distribution family and a novel approach for converting deterministic weights to probabilistic distributions without training. The complete algorithmic implementation is provided in Sec. 3.4, with theoretical foundations addressed in a separate section (Sec. 4).
Notation. Scalars, vectors, and matrices are denoted by lowercase letters, lowercase boldface letters, and uppercase boldface letters, respectively. For a matrix ${\bm{X}}=[{\bm{x}}_{1},\cdots,{\bm{x}}_{n}]\in\mathbb{R}^{m\times n}$, we use $\operatorname{vec}({\bm{X}})=[{\bm{x}}_{1}^{\top},{\bm{x}}_{2}^{\top},\cdots,{\bm{x}}_{n}^{\top}]^{\top}\in\mathbb{R}^{(mn)\times 1}$ to denote vectorization. $\otimes$ and $\circ$ denote the Kronecker and element-wise product, respectively. We use $\mathbf{0}_{n}\in\mathbb{R}^{n\times n}$ to denote a zero matrix.
@@ 3.1 Preliminaries
Low-Rank Adaptation (LoRA). Given a pre-trained neural network layer with weight matrix ${\bm{W}}_{0}$, Low-Rank Adaptation (LoRA) [23] confines weight updates to a low-rank subspace during fine-tuning, expressing the update as $\Delta{\bm{W}}={\bm{B}}{\bm{A}}$, where $\Delta{\bm{W}}\in\mathbb{R}^{m\times n}$, ${\bm{B}}\in\mathbb{R}^{m\times r}$, and ${\bm{A}}\in\mathbb{R}^{r\times n}$. For input ${\bm{h}}$ and output ${\bm{z}}$ of the LoRA layer, the forward pass computation is then given by:
@@
LoRA Bayesianization with Low-Rank Gaussian Distribution. BLoB [61], a pioneering work in low-rank Bayesianization for LLMs, empirically demonstrates that modeling ${\bm{A}}$’s elements with independent Gaussian variables suffices for effective uncertainty estimation in LoRA. Specifically, the probability density of each element of ${\bm{A}}$ follows $q(A_{ij})={\mathcal{N}}(A_{ij}|M_{ij},\Omega_{ij}^{2}),\forall i\in[r],\forall j\in[n]$, where matrices ${\bm{M}}$ and ${\bm{\Omega}}$, sharing the dimensions of ${\bm{A}}$, represent the mean and standard deviation of the random variable ${\bm{A}}$, respectively. This formulation is equivalent to approximating the Bayesianized low-rank adapter’s posterior in the full-weight space of ${\bm{W}}$ with a low-rank degenerate distribution:
@@
where ${\bm{\theta}}=\{{\bm{M}},{\bm{\Omega}}\}$ denotes the set of parameters for modeling ${\bm{A}}$’s posterior distribution, ${\bm{\mu}}_{q}=\operatorname{vec}({\bm{W}}_{0}+{\bm{B}}{\bm{M}})$ is its mean, and ${\bm{\Sigma}}_{q}=[{\bm{I}}_{n}\otimes{\bm{B}}][\operatorname{diag}(\operatorname{vec}({\bm{\Omega}})^{2})][{\bm{I}}_{n}\otimes{\bm{B}}^{\top}]$ is its low-rank degenerate covariance matrix. In this paper, we adopt a similar approach for modeling the variational distribution of the weight posterior, focusing exclusively on Bayesianizing the weight update matrix ${\bm{A}}$.
3.2 TFB’s Variational Low-Rank Isotropic Gaussians
@@
where ${\bm{\mu}}_{q}$ is defined as in Eqn. 2. Here, $\sigma_{q}^{2}{\bm{I}}\in\mathbb{R}^{mn\times mn}$ represents a full-rank isotropic covariance matrix with standard deviation $\sigma_{q}$, and $\operatorname{proj}(\cdot)$ denotes a linear projection operator that maps the full-space covariance matrix onto the low-rank space (see Proposition D.1.1 for details). 111 $\operatorname{proj}(\cdot)$ only depends on the rank $r$ of the trained LoRA.
@@
TFB as Generalized Variational Inference. The choice of low-rank isotropic Gaussian approximate posteriors serves both theoretical and empirical purposes: it provides a single-parameter family that enables converting the generalized variational inference into a variance maximization problem (more details in Sec. 3.3, Theorem 4.2, and Appendix F.1), and empirically outperforms alternative distribution families (Sec. 5.3). Below, we present a practically efficient implementation for Bayesianizing LoRA under the constraint specified in Eqn. 3, with detailed theoretical analysis provided in Theorem 4.1.
@@
TFB in Practice. Consider a LoRA layer with weight updates ${\bm{B}}\in\mathbb{R}^{m\times r},{\bm{A}}\in\mathbb{R}^{r\times n}$ and a standard deviation scale $\sigma_{q}>0$. We begin by computing the compact Singular Value Decomposition (SVD) [33] of ${\bm{B}}$:
${\bm{B}}$ $={\bm{U}}\operatorname{diag}({\bm{d}}){\bm{V}}^{\top},$ (4)
@@
where ${\bm{U}}\in\mathbb{R}^{m\times r}$ and ${\bm{V}}\in\mathbb{R}^{r\times r}$ are orthonormal matrices, and ${\bm{d}}=[d_{1},d_{2},\cdots,d_{r}]^{\top}$ is the vector consisting of singular values with all positive entries222 By stating ${\bm{d}}\succ{\bm{0}}$, we assume ${\bm{B}}$ has the full column rank $r$, which usually holds for LLM adaptation. . We then transform the original weight matrices $\{{\bm{B}},{\bm{A}}\}$ into an equivalent pair
$\{{\bm{B}}^{\prime}$ $={\bm{U}}\operatorname{diag}({\bm{d}}),{\bm{A}}^{\prime}={\bm{V}}^{\top}{\bm{A}}\},$ (5)
@@
maintaining the equality $\Delta{\bm{W}}={\bm{B}}{\bm{A}}={\bm{B}}^{\prime}{\bm{A}}^{\prime}$. Following BLoB’s Asymmetric Bayesianization scheme, we define the variational distribution for ${\bm{A}}^{\prime}$ using the mean matrix ${\bm{M}}={\bm{A}}^{\prime}$ and the standard deviation matrix ${\bm{\Omega}}\in\mathbb{R}^{r\times n}$, such that
$q(A^{\prime}_{ij})={\mathcal{N}}(A^{\prime}_{ij}|M_{ij},\Omega_{ij}^{2}),\forall i\in[r],\forall j\in[n].$ (6)
Unlike BLoB, our ${\bm{\Omega}}$ is not freely parameterized but instead derived from projecting the full-space matrix $\sigma_{q}{\bm{I}}$ onto the low-rank weight space:
$\Omega_{ij}$ $=\nicefrac{{\sigma_{q}}}{{d_{i}}},\quad\forall i\in[r],\forall j\in[n],$ (7)
where ${\bm{d}}$ is defined in Eqn. 4. This solution can be expressed compactly as ${\bm{\Omega}}=[\nicefrac{{\sigma_{q}}}{{{\bm{d}}}},\cdots,\nicefrac{{\sigma_{q}}}{{{\bm{d}}}}]$, comprising $n$ repeated vectors. To summarize, our TFB
  • takes as input a trained LoRA matrix pair $\{{\bm{B}}={\bm{U}}\operatorname{diag}({\bm{d}}){\bm{V}}^{\top},{\bm{A}}\}$ and a predetermined standard deviation $\sigma_{q}$, and
  • outputs a “Bayesianized” LoRA adapter $\{{\bm{B}}^{\prime},{\bm{A}}^{\prime}\}$, where ${\bm{B}}^{\prime}={\bm{U}}\operatorname{diag}({\bm{d}})$, and ${\bm{A}}^{\prime}$ becomes a distribution $q({\bm{A}}^{\prime})=\prod_{i\in[r],j\in[n]}{\mathcal{N}}(A^{\prime}_{ij}|M_{ij},\Omega_{ij}^{2})$, with ${\bm{M}}={\bm{V}}^{\top}{\bm{A}}$, and ${\bm{\Omega}}=[\nicefrac{{\sigma_{q}}}{{{\bm{d}}}},\cdots,\nicefrac{{\sigma_{q}}}{{{\bm{d}}}}].$
@@
Note that the formulation in Eqn. 7 significantly improves memory efficiency during inference, reducing the storage for standard deviation parameters from $O(rn)$ to $O(r)$. While alternative parameterization approaches are possible, they must be capable of generating the low-rank isotropic Gaussian noises as demonstrated in Theorem 4.1. We have selected the current method (implementation) to ensure maximum compatibility with existing codebases [61]. In TFB, we use a single $\sigma_{q}$ shared across all LoRA layers.
3.3 TFB as Variance Maximization